OpenStudy (leeanm):

The coordinates of the vertices of a rectangle are (−3, 4), (7, 2) , (6, −3), and (−4, −1) . What is the perimeter of the rectangle to the nearest tenth of a unit? Enter your answer as a decimal in the box. Round each step to the nearest tenth.

563blackghost (563blackghost):

Use distance formula... \(\huge\bf{d=\sqrt{(y_{2}-y_{1})^{2}+(x_{2}-x_{1})^{2}}}\) Find the distance between `(-3,4) and (7,2) then multiply by 2.` Then find the distance between `(7,2) and (6,-3) then multiply by 2.` Once this is done you would add the two distances and find your answer.

OpenStudy (leeanm):

how do i do that?

563blackghost (563blackghost):

Plug in the points... \(\huge\bf{d=\sqrt{(2-4)^{2}+(7-(-3))^{2}}}\)

OpenStudy (leeanm):

i dont know how to do that.... im not very good at math..

563blackghost (563blackghost):

We need to simplify...so we would subtract the parenthesis... \(\huge\bf{d=\sqrt{(-2)^{2}+(10)^{2}}}\) Next we square them... \(\huge\bf{d=\sqrt{4+100}}\) Add.. \(\huge\bf{d=\sqrt{104}=2\sqrt{26}}\) We would multiply this by 2... \(\huge\bf{(2\sqrt{26})\times 2 = 4\sqrt{26}}\) This is the distance total of two opposite sides...

563blackghost (563blackghost):

Now we find the next two points :) \(\huge\bf{d=\sqrt{(-3 - 2)^{2}+(6-7)^{2}}}\) Subtract... \(\huge\bf{d=\sqrt{(-5)^{2}+(1)^{2}}}\) Square... \(\huge\bf{d= \sqrt{25 + 1}}\) Add.. \(\huge\bf{d= \sqrt{26}}\) In which we would multiply by 2.... \(\huge{(\sqrt{26}) \times 2 = 2\sqrt{26}}\)

563blackghost (563blackghost):

You would then add the two.... \(\huge\bf{2 \sqrt{26} + 4 \sqrt{26} = ?}\)

563blackghost (563blackghost):

@leeanm

563blackghost (563blackghost):

Plug this in to your calculator....

OpenStudy (leeanm):

30.5?

563blackghost (563blackghost):

Close. If you plug into your cal you would get `30.59411708` you would need to round to the nearest tenth so it would be `30.6`

OpenStudy (leeanm):

so 30.6 is my answer?

563blackghost (563blackghost):

Yes

OpenStudy (leeanm):

ok!

OpenStudy (leeanm):

Can you help me with another?

563blackghost (563blackghost):

I can't. Sorry I have to go though I think @Will.H can he's one great owlet :)

OpenStudy (leeanm):

ok

563blackghost (563blackghost):

But post it in a new post to not crowd this one....Oh and... \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

OpenStudy (will.h):

oh amazing work @563blackghost

563blackghost (563blackghost):

Why thank you @Will.H :)

OpenStudy (will.h):

well that's a really remarkable work! great job :)

OpenStudy (will.h):

@leeanm if you have other question post them up in a different post and tag me like @Will.H